jagr2713 (jagr2713):

Harder Trick Question: From a book, a number of consecutive pages are missing. The sum of the page numbers of these pages is 9808. Which pages are missing?

jagr2713 (jagr2713):

@pooja195 let someone else try this

jagr2713 (jagr2713):

@rational @Miracrown @Zale101 @KingGeorge @sleepyjess @Nnesha

pooja195 (pooja195):

i dont know this one ;-;

jagr2713 (jagr2713):

oh lol lets see if someone else knows

jagr2713 (jagr2713):

There is a trick word in there :DD

jagr2713 (jagr2713):

Hint: pages (plural)

jagr2713 (jagr2713):

You guys got this :D

jagr2713 (jagr2713):

Hint: Let the number of missing pages be n and the first missing page p+1.

OpenStudy (anonymous):

pages 291 thru 322 are missing!!!! Am i right?

jagr2713 (jagr2713):

Explain

jagr2713 (jagr2713):

*Waiting lol*

jagr2713 (jagr2713):

you guys give up?

pooja195 (pooja195):

no

jagr2713 (jagr2713):

Ok .*Jeopardy Song*

OpenStudy (anonymous):

Sorry my internet keeps restarting and everytime it does my explanation goes away...

jagr2713 (jagr2713):

oh ok

OpenStudy (sleepyjess):

I say that 2 through 9806 are missing

jagr2713 (jagr2713):

lol @sleepyjess but no lol u gave up yet @pooja195

OpenStudy (sleepyjess):

Why am I wrong though? :(

jagr2713 (jagr2713):

ull see lets wait for POOJA

OpenStudy (sleepyjess):

Oooohhhh nvm

jagr2713 (jagr2713):

Ok @pooja195 i am going to answer it in the next 30seconds lol

pooja195 (pooja195):

number of missing pages be n and the first page p+1. Then the pages p+1 up to and including p+n are missing, and n times the average of the numbers of the missing pages must be equal to 9808: n×( ((p+1)+(p+n))/2 ) = 9808 In other words: n×(2×p+n+1)/2 = 2×2×2×2×613 So: n×(2×p+n+1) = 2×2×2×2×2×613 One of the two terms n and 2×p+n+1 must be even, and the other one must be odd. Moreover, the term n must be smaller than the term 2×p+n+1. It follows that there are only two solutions: n=1 and 2×p+n+1=2×2×2×2×2×613, so n=1 and p=9808, so only page 9808 is missing. n=2×2×2×2×2 and 2×p+n+1=613, so n=32 and p=290, so the pages 291 up to and including 322 are missing. Because it is asked which pages are missing, the solution is: the pages 291 up to and including 322 are missing.

pooja195 (pooja195):

Refresh the page -_-

jagr2713 (jagr2713):

FINALLY YESSSSS

pooja195 (pooja195):

:P this is why i said to wait

OpenStudy (sleepyjess):

O_O

jagr2713 (jagr2713):

SEE why @sleepyjess

pooja195 (pooja195):

WHY DID IT HAVE TO BE LONG XD

jagr2713 (jagr2713):

Lol but great job :D @KingGeorge

OpenStudy (kinggeorge):

What did I do here lol

pooja195 (pooja195):

xD

jagr2713 (jagr2713):

nah just showing you lol

OpenStudy (sleepyjess):

lol

jagr2713 (jagr2713):

Well i am done for the day ~Thanks for playing guys, Have a great night :D